In an experiment, we have isolated 0.25g of caffeine from 3 tea bags. Please calculate the number of tea bags needed if we want to obtain 345mmol of caffeine.

Please show your work, thank you!
Answer #1

345mmol of caffeine.    = 0.345 moles of of caffeine.

gram molar mass of caffeine.    = 194.19g/mole

mass of caffeine.      = no of moles * gram molar mass

                                 = 0.345*194.19

                                 = 67g

0.25g of caffeine   = 3 tea bags

67g of caffeine = 3*67/0.25   = 804 tea bags

804 tea bags >>>>answer